OpenStudy (anonymous):

3(x-2) +2 = 5x -7 -2x

OpenStudy (ajprincess):

\( 3(x-2) +2 = 5x -7 -2x\) \(3*x-3*2+2=5x-7-2x\) \(3x-6+2=5x-7-2x\) \(3x-4=3x-7\) See if u can do it nw.

OpenStudy (jiteshmeghwal9):

3x-6+2=5x-7-2x 3x-4=3x-7 equation do not exist

OpenStudy (ajprincess):

ya u right @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

thanx @ajprincess :)

OpenStudy (ajprincess):

yw:)

OpenStudy (anonymous):

thank you! test is today and this one was on the study guide. I have one more. . . 5z = -3(z +7)

OpenStudy (ajprincess):

5z=-3(z+7) 5z=-3*z-3*7 5z=-3z-21 See if u can do it nw